# A Compact Implicit Neural Representation for Efficient Storage of Massive 4D Functional Magnetic Resonance Imaging

### 30 Nov 2023 v1

@@ Abstract
Functional Magnetic Resonance Imaging (fMRI) data is a kind of widely used four-dimensional biomedical data, demanding effective compression but presenting unique challenges for compression due to its intricate temporal dynamics, low signal-to-noise ratio, and complicated underlying redundancies. This paper reports a novel compression paradigm specifically tailored for fMRI data based on Implicit Neural Representation (INR). The proposed approach focuses on removing the various redundancies among the time series, including (i) conducting spatial correlation modeling for intra-region dynamics, (ii) decomposing reusable neuronal activation patterns, and using proper initialization together with nonlinear fusion to describe the inter-region similarity. The above scheme properly incorporates the unique features of fMRI data, and experimental results on publicly available datasets demonstrate the effectiveness of the proposed method, surpassing state-of-the-art algorithms in both conventional image quality evaluation metrics and fMRI downstream tasks. This work in this paper paves the way for sharing massive fMRI data at low bandwidth and high fidelity.
Figure 1: The trace and slice comparison of an exemplar fMRI data modeled by previous INR-based methods (NeRF and NeRV) and the proposed approach.
@@ 1 Introduction
Functional Magnetic Resonance Imaging (fMRI), as a widely available non-invasive imaging tool, has been widely used in cognitive neuroscience, clinical psychology, and psychiatry. As large-scale fMRI datasets continue to proliferate, coupled with the ever-growing demand in both clinical [53, 10, 38] and academic domains [15, 24, 40], the fMRI data constitutes a paramount proportion of massive biomedical data [51, 19, 12, 29]. For efficient storage and transmission, there is an urgent necessity for a high-quality data compression paradigm tailored specifically to fMRI.
@@
fMRI data encompasses three spatial dimensions in addition to a temporal dimension, with the temporal dimension presenting intricate neural dynamics [23] and with a relatively low signal-to-noise ratio [11], which pose significant challenges for conventional biomedical data compression algorithms. Moreover, distinct from natural videos or ultrasound data, the redundancies in fMRI data predominantly reside within and between temporal signals of various brain regions [32, 52], rather than adjacent frames. This unique characteristic also challenges conventional compression algorithms and calls for new compressors that make extensive use of the intrinsic structure of fMRI data.
@@
Implicit Neural Representation (INR) is becoming increasingly widespread in various domains, such as shape representation [18, 17, 31], scene rendering [33, 62, 57, 61], and image/video representation [47, 7, 6, 5]. Due to its inherent advantages in modeling internal data correlations [46], as well as the recent development of INR-based data compression algorithms [13, 6, 49, 14, 60, 58, 59] achieving comparable or even surpassing traditional compressors, INR stands out as the most cutting-edge and promising approach in the fields of data compression and deep learning. The success of INR-based compression on videos and three-dimensional biomedical data implies the feasibility but still cannot directly support its applications on fMRI data. First, it is straightforward but difficult to directly extend existing biomedical data compression techniques [61] to 4D fMRI data. When directly employing INR to model the mapping from spatio-temporal coordinates to the signal values, the intricate dynamics [23] and heavy noise [11] might degenerate the coding accuracy and result in poor encoding accuracy in the temporal dimension, as shown in Fig. 1. Second, the INR-based compression algorithms designed for conventional videos, such as NeRV, [6], focuses on eliminating redundancies in adjacent video frames, and is not applicable to fMRI structure. The comparison is shown in Fig. 1. In summary, one should design new compression techniques to incorporate the unique features of fMRI data, including the structure, redundancy, and imaging quality.
@@
To achieve high-quality compression of 4D fMRI data, we propose an INR-based paradigm, leveraging INR’s powerful representation capability and the unique features of the fMRI data. First, we employ a neural network representing the coordinates-to-intensity mapping to model the correlations of neural dynamics within the same brain region, thereby eliminating redundancies. Next, we decompose the data into a set of reusable neuronal activation patterns and corresponding spatial distribution, effectively removing the among-region redundancies. Subsequently, we use INR to reduce redundancies underlying the spatial distribution of activation patterns and introduce a feature fusion network to simulate the integration of neuronal activation patterns in real fMRI data, thus enhancing the encoding precision. Finally, we introduce an initialization scheme for neuronal activation patterns based on Independent Component Analysis (ICA) to further decrease correlations between different neuronal activation patterns.
@@
We demonstrate the largely superior performance on publicly available datasets to validate the proposed method. It’s worth noting that, in terms of evaluation metrics, in addition to the common metrics like PSNR and SSIM, which measure pixel-wise differences, we also introduced additional evaluation criteria widely used in fMRI data analysis as part of our assessment. Because fMRI data is primarily utilized for data analysis rather than diagnosis by observation, using the disparity between decompressed data and the original version in downstream tasks as an additional evaluation metric allows for a more comprehensive assessment of the compression algorithms.
@@
  • Superb Compression Performance. We achieve advantageous data fidelity at high compression ratios, in terms of both image quality and downstream tasks.
Figure 2: The workflow and the basic structure of the proposed fMRI compression approach.
@@ 2 Related Work
Data compression plays a significant role in the storage and sharing of data. Over the past few decades, there has been rapid development in data compression algorithms, leading to tremendous successes with algorithms like JPEG[55], H.264[56], and H.265[50]. The techniques which have been adopted by these compressors such as discrete wavelet transform[22] and block-size motion compensation later became extensively applied in the field of data compression. In recent years, with the emergence of deep learning, data compression techniques based on deep learning have gained momentum, resulting in the development of various new compression algorithms[30, 1]. However, when it comes to high-dimensional data, such as four-dimensional data, there still lack proper compressors.
INR-based Compression.
@@
Recent progress in data compression has introduced INR for more compact data representation, yielding promising results[6, 27, 13, 8]. Unlike traditional data compression algorithms that employ explicit encoding methods, INR-based data compression algorithms leverage the powerful information capacity of neural networks to implicitly store data information within network parameters, thereby achieving efficient compression. However, for the task of fMRI compression, existing INR-based data compression algorithms have certain limitations. Some algorithms focus only on spatial or temporal redundancies, making it difficult to extend them to four-dimensional datasets or achieve satisfactory compression results. Some INR-based video compression algorithms address both types of redundancies, but due to the higher demands for fidelity in medical image compression compared to natural images, and the increased complexity of fMRI data structures, the compression strategies employed by existing INR compression algorithms are challenging to apply. In order to more effectively preserve fMRI information, we have altered the data modeling approach of INR. Instead of directly modeling fMRI signals, we utilize INR to model the interrelationships among brain region activities.
High Dimension Medical Data Compression.
@@
There is a lack of adequate compression methods for four-dimensional medical data. The technical routes of these algorithms can be roughly divided into two categories. The first category is based on motion compensation, which use motion vectors to model the difference between frames to reduce redundancies[37, 45, 43, 44]. However, unlike nature scene videos, the redundancies in fMRI data is concentrated between the temporal signals within and between various brain regions, rather than between adjacent frames[32, 52]. Therefore, using motion compensation does not effectively utilize the spatiotemporal redundancies of fMRI for compression. The second category is based on transform like wavelet transform.[25, 28, 39] The transform applied to images or signals is modified to higher dimensions and then applied to medical imaging data. This modification is usually a combination of 1D transform and 3D transform, leading to limited ability to sparsely capture more complex, higher-order discontinuities[4]. Our method is completely different from the traditional data compression algorithms in terms of spatiotemporal redundancies compression. We use the correlations between brain region activities in fMRI to perform redundancies compression, fully considering the characteristics of the data.
@@ 3.1 Mathematical Representation of fMRI data
$\begin{split}&\mathbf{X}=\{\boldsymbol{x}_{\boldsymbol{v}}|\boldsymbol{v}^{(0)}\in[0,W),\\ &\boldsymbol{v}^{(1)}\in[0,H),\boldsymbol{v}^{(2)}\in[0,D)\}.\end{split}$ (1)
@@
The principle of localization in brain function organization suggests that brain functions are carried out in a set of brain regions [32]. This implies that the brain space can be divided into several brain regions based on function, with similarities in neuronal activation within each brain region. Therefore, the set of time series for each brain region can be represented as
@@
in which $\mathbf{V}_{i}$ represents the sets of three-dimensional coordinates for the brain region labeled as $i$, and $N$ represents the number of brain region labels. Therefore, we have
@@ 3.2 Correlation Modeling with Spatial Coordinates
The overall structure of the network is shown in Fig. 2. Combined with the mathematical representation of fMRI, our neural network parameterization function is designed at a macro level as
$f(\boldsymbol{v}|\theta)\rightarrow\boldsymbol{x}_{\boldsymbol{v}}.$ (5)
Here $\theta$ represents the network parameters. The high similarities within the neural activation in a brain region mentioned above indicates that there exist high correlations between $\boldsymbol{x}_{\boldsymbol{v}}$ in the same $\mathbf{V}_{i}$, which matches with $f$’s advantageous modeling capability of local structures. In other words, if we use the neural network parameterization function mapping from spatial coordinates $\boldsymbol{v}$ to the intensity, the local correlations of the function $f$ will be enhanced. This aligns well with the advantage of INR in modeling internal correlations, and can more effectively eliminate the redundancies of time-series within local brain regions.
3.3 Neuronal Activation Patterns Decomposition
The connectionism principle of brain function suggests that there exist correlations within the neuronal activation patterns in certain brain regions[32, 52]. At the same time, theoretically, the generation process of fMRI signals can be viewed as the output of a Linear Time-Invariant System (LTI System)[2, 3]:
$\boldsymbol{x}_{\boldsymbol{v}}^{(t)}=H(\sum_{i=0}^{K-1}\boldsymbol{a}_{\boldsymbol{v}}^{(i)}\cdot u_{i}(t))=\sum_{i=0}^{K-1}\boldsymbol{a}_{\boldsymbol{v}}^{(i)}\cdot H(u_{i}(t)).$ (6)
Here, $H$ denotes the system generating fMRI signals, $u_{i}(t)$ is the input stimulus, $\boldsymbol{a}_{\boldsymbol{v}}^{(i)}$ is the stimulus distribution of signal intensity, and $H(u_{i}(t))$ is the neuronal activation pattern. Therefore, theoretically, $\boldsymbol{x}_{\boldsymbol{v}}$ can be decomposed into the weighted superposition of several time series characterizing neuronal activation patterns. Inspired by this prior, as shown in Fig. 2, we established a learnable matrix $C_{\text{train}}\in\mathbb{R}^{K\times T}$. Each row of the matrix is mapped to the feature space through Conv1d blocks to obtain the feature matrix $F$. As such, the network can learn $K$ types of reusable neuronal activation patterns, and the correlations of neuronal activation patterns between brain regions can be modeled by the neural network, thereby eliminating the redundancies of time series between brain regions.
@@ 3.4 Modeling of Activation Pattern Distribution and Feature Fusion
While modeling several reusable neuronal activation patterns, we use INR groups to model their spatial distribution
$g_{i}(\boldsymbol{v}|\sigma_{i})\rightarrow\boldsymbol{w}^{(i)},i\in[0,K),$ (7)
and then concatenate the outputs into
$\boldsymbol{w}=[\boldsymbol{w}^{(0)},\boldsymbol{w}^{(1)},\cdots,\boldsymbol{w}^{(K-1)}].$ (8)
Here function $g_{i}$ represents the $i$th INR network with $\sigma_{i}$ being its parameters. The components of $\boldsymbol{w}\in\mathbb{R}^{K}$ represent the spatial distribution of each activation pattern. In this way, we can further eliminate the redundancies within the distribution of neuronal activities.
@@ 3.5 Activation Pattern Initialization by ICA
To enable the network to converge quickly, it is necessary to set good initialization to the learnable matrix $C_{train}$. We initialize the matrix with a set of independent values to minimize the redundancies within each activation pattern. Specifically, we adopted the ICA algorithm, decomposing the original fMRI data into a set of time-series bases. The matrix form of ICA can be represented as $Y=AS$, where $Y$ is the original signal matrix and decomposed into a mixing matrix $A$ and a source matrix $S=\{\vec{s}_{1},\vec{s}_{2},\cdots\}$. Correspondingly, applying ICA to fMRI signals gets
$\boldsymbol{x}_{\boldsymbol{v}}=\boldsymbol{a}_{\boldsymbol{v}}\cdot\left[\boldsymbol{c}_{0},\boldsymbol{c}_{1},\cdots,\boldsymbol{c}_{K-1}\right]^{T}.$ (9)
In this equation, $\boldsymbol{x}_{\boldsymbol{v}}$ represents the fMRI time series, $\boldsymbol{c}_{i}\in\mathbb{R}^{T\times 1}$ denotes the activation patterns with the vector $\boldsymbol{a}_{\boldsymbol{v}}=[\boldsymbol{a}_{\boldsymbol{v}}^{(0)},\boldsymbol{a}_{\boldsymbol{v}}^{(1)},\cdots,\boldsymbol{a}_{\boldsymbol{v}}^{(K-1)}]$ describing their spatial distributions. These activation patterns will be used for the initialization of the matrix $C_{train}$ to minimize the dependency among matrix rows.

### 29 Feb 2024 v2

@@ Abstract
Functional Magnetic Resonance Imaging (fMRI) data is a widely used kind of four-dimensional biomedical data, which requires effective compression. However, fMRI compressing poses unique challenges due to its intricate temporal dynamics, low signal-to-noise ratio, and complicated underlying redundancies. This paper reports a novel compression paradigm specifically tailored for fMRI data based on Implicit Neural Representation (INR). The proposed approach focuses on removing the various redundancies among the time series by employing several methods, including (i) conducting spatial correlation modeling for intra-region dynamics, (ii) decomposing reusable neuronal activation patterns, and (iii) using proper initialization together with nonlinear fusion to describe the inter-region similarity. This scheme appropriately incorporates the unique features of fMRI data, and experimental results on publicly available datasets demonstrate the effectiveness of the proposed method, surpassing state-of-the-art algorithms in both conventional image quality evaluation metrics and fMRI downstream tasks. This work in this paper paves the way for sharing massive fMRI data at low bandwidth and high fidelity.
Keywords:
fMRI INR Compression
@@ 1 Introduction
Functional Magnetic Resonance Imaging (fMRI), as a widely available non-invasive imaging tool, has been widely used in cognitive neuroscience, clinical psychology, and psychiatry. As large-scale fMRI datasets continue to proliferate, coupled with the ever-growing demand in both clinical [53, 10, 38] and academic domains [15, 24, 40], the fMRI data constitutes a significant proportion of massive biomedical data [51, 19, 12, 29]. For efficient storage and transmission, there is an urgent necessity for a high-quality data compression paradigm tailored specifically to fMRI.
@@
fMRI data consists of three spatial dimensions in addition to a temporal dimension, with the temporal dimension presenting intricate neural dynamics [23] but having a relatively low signal-to-noise ratio [11], which poses significant challenges for conventional biomedical data compression algorithms. Moreover, distinct from natural videos or ultrasound data, the redundancies in fMRI data predominantly reside within and between temporal signals of various brain regions [32, 52], rather than adjacent frames. This unique characteristic also challenges conventional compression algorithms and calls for new compressors that make extensive use of the intrinsic structure of fMRI data.
@@
Implicit Neural Representation (INR) is becoming increasingly widespread in various domains, such as shape representation [18, 17, 31], scene rendering [33, 62, 57, 61], and image/video representation [47, 7, 6, 5]. Due to its inherent advantages in modeling internal data correlations [46], the recent advancements in INR-based data compression algorithms [13, 6, 49, 14, 59, 58, 60] can rival or even surpass traditional compressors, and INR stands out as the most cutting-edge and promising approach in the fields of data compression and deep learning. The success of INR-based compression on videos and three-dimensional biomedical data implies the feasibility but still cannot directly support its application to fMRI data. First, it is straightforward but difficult to directly extend existing biomedical data compression techniques [61] to 4D fMRI data. When directly employing INR to model the mapping from spatio-temporal coordinates to the signal values, the intricate dynamics [23] and heavy noise [11] might result in poor encoding accuracy in the temporal dimension, as shown in Fig. 1. Second, the INR-based compression algorithms designed for conventional videos, such as HNeRV, [5], focus on eliminating redundancies in adjacent video frames, and is not applicable to fMRI structure. The comparison is shown in Fig. 1. In summary, new compression techniques need to be designed to incorporate the unique features of fMRI data, including the structure, redundancy, and imaging quality.
Figure 1: The trace and slice comparison of an exemplar fMRI data modeled by previous INR-based methods (NeRF and HNeRV) and the proposed approach.
@@
To achieve high-quality compression of 4D fMRI data, we propose an INR-based paradigm, leveraging INR’s powerful representation capability and the unique features of the fMRI data. First, we employ a neural network to model the coordinates-to-intensity mapping, capturing the correlations of neural dynamics within the same brain region, thereby eliminating redundancies. Next, we decompose the data into a set of reusable neuronal activation patterns and their corresponding spatial distribution, effectively removing the among-region redundancies. Subsequently, we use INR to reduce redundancies in the spatial distribution of activation patterns and introduce a feature fusion network to simulate the integration of neuronal activation patterns in real fMRI data, enhancing the encoding precision. Finally, we introduce an initialization scheme for neuronal activation patterns based on Independent Component Analysis (ICA) to find the primary neuronal activation patterns among the fMRI data, which can better retain the more significant activation patterns across brain regions to align the requirement of fidelity.
@@
We demonstrate the largely superior performance of our method on publicly available datasets in order to validate it. It’s worth noting that, in terms of evaluation metrics, in addition to the common metrics like PSNR and SSIM, which measure pixel-wise differences, we also introduced additional evaluation criteria as part of our assessment which are downstream tasks widely used in fMRI data analysis. Given that fMRI data is primarily utilized for data analysis rather than diagnosis observation, using the disparity between the decompressed data and the original version in downstream tasks as an additional evaluation metric allows for a more comprehensive assessment of the compression algorithms.
@@
  • Initialization with Main Neuronal Activation Patterns by ICA. An ICA-based method to identify the dominant activation patterns in brain regions for initialization, in order to preserve salient information better.
Figure 2: The workflow and the basic structure of the proposed compression approach.
@@ 2 Related Work
Data Compression. Data compression plays a significant role in the storage and sharing of data. Over the past few decades, there has been rapid development in data compression algorithms, leading to tremendous successes with algorithms like JPEG[55], H.264[56], and H.265[50]. The techniques which have been adopted by these compressors such as discrete wavelet transform[22] and block-size motion compensation,have later become extensively applied in the field of data compression. In recent years, with the emergence of deep learning, data compression techniques based on deep learning have gained momentum, resulting in the development of various new compression algorithms[30, 1]. However, when it comes to high-dimensional data, such as four-dimensional data, proper compressors are still lacking.
@@
INR-based Compression. Recent progress in data compression has introduced INR for more compact data representation, yielding promising results[6, 27, 13, 8]. Unlike traditional data compression algorithms that employ explicit encoding methods, INR-based data compression algorithms leverage the powerful information capacity of neural networks to implicitly store data information within network parameters, thereby achieving efficient compression. However, for the task of fMRI compression, existing INR-based data compression algorithms have certain limitations. Some algorithms focus only on spatial or temporal redundancies, making it difficult to extend them to four-dimensional datasets or achieve satisfactory compression results. Some INR-based video compression algorithms address both types of redundancies, but due to the higher demands for fidelity in medical image compression compared to natural images, and the increased complexity of fMRI data structures, the compression strategies employed by existing INR compression algorithms are challenging to apply. In order to more effectively preserve fMRI information, we have altered the data modeling approach of INR. Instead of directly modeling fMRI signals, we utilize INR to model the interrelationships among brain region activities.
@@
High Dimension Medical Data Compression. There is a lack of adequate compression methods for four-dimensional medical data. The technical routes of these algorithms can be roughly divided into two categories. The first category is based on motion compensation, which uses motion vectors to model the difference between frames to reduce redundancies[37, 45, 43, 44]. However, unlike nature scene videos, the redundancies in fMRI data are concentrated between the temporal signals within and between various brain regions, rather than between adjacent frames[32, 52]. Therefore, using motion compensation does not effectively utilize the spatiotemporal redundancies of fMRI for compression. The second category is based on transform ,such as wavelet transform and so on.[25, 28, 39] The transform applied to images or signals is modified to higher dimensions and then applied to medical imaging data. This modification is usually a combination of 1D transform and 3D transform, leading to limited ability to sparsely capture more complex, higher-order discontinuities[4]. Our method is completely different from the traditional data compression algorithms in terms of spatiotemporal redundancies compression. We leverage the correlations between brain region activities in fMRI to perform redundancies compression, fully considering the characteristics of the data.
@@ 3.1 Mathematical Representation of fMRI data
$\mathbf{X}=\{\boldsymbol{x}_{\boldsymbol{v}}|\boldsymbol{v}^{(0)}\in[0,W),\boldsymbol{v}^{(1)}\in[0,H),\boldsymbol{v}^{(2)}\in[0,D)\}.$ (1)
@@
The principle of localization in brain function organization suggests that brain functions are carried out in a set of brain regions [32]. This implies that the brain space can be divided into several brain regions based on function, and there are similarities in neuronal activation within each brain region. Therefore, in fMRI data, the set of time series for each brain region can be represented as
@@
in which $\mathbf{V}_{i}$ represents the set of three-dimensional coordinates for the brain region labeled as $i$, and $N$ represents the number of brain region labels. Therefore, we have
@@ 3.2 Correlation Modeling with Spatial Coordinates
The high similarities within the neural activation in a brain region mentioned above indicate that there are high correlations between $\boldsymbol{x}_{\boldsymbol{v}}$ in the same $\mathbf{V}_{i}$, and the region $\mathbf{V}_{i}$ is usually a set of continuous space in brain. In other words, if we model the fMRI data as the mapping from spatial coordinates $\boldsymbol{v}$ to the signal time series $\boldsymbol{x}_{\boldsymbol{v}}$
$F(\boldsymbol{v})\rightarrow\boldsymbol{x}_{\boldsymbol{v}},$ (5)
the local correlations of the function $F$ will be considerable, which matches with the advantages of INR-based methods that has superior capability of local structure modeling. Integrating the mathematical representation of fMRI, our parameterization function for neural network is designed at a macro level as
$f(\boldsymbol{v}|\theta)\rightarrow\widetilde{\boldsymbol{x}_{\boldsymbol{v}}}.$ (6)
Here $\theta$ represents the network parameters, and $\widetilde{\boldsymbol{x}_{\boldsymbol{v}}}$ represents the prediction of the ground truth $\boldsymbol{x}_{\boldsymbol{v}}$. The overall structure of the network is shown in Fig. 2. This design aligns well with the strength of INR in modeling internal correlations, and can more effectively eliminate the redundancies of time series within local brain regions.
3.3 fMRI time-series signal Decomposition
The connectionism principle of brain function suggests that there exist correlations within the neuronal activation patterns in certain brain regions[32, 52]. Simultaneously, the generation process of fMRI time-series signals can be theoretically viewed as the output of a Linear Time-Invariant System (LTI System)[2, 3]:
$\boldsymbol{x}_{\boldsymbol{v}}^{(t)}=H(\sum_{i=0}^{K-1}\boldsymbol{w}_{\boldsymbol{v}}^{(i)}\cdot u_{i}(t))=\sum_{i=0}^{K-1}\boldsymbol{w}_{\boldsymbol{v}}^{(i)}\cdot H(u_{i}(t)).$ (7)
Here, $H$ denotes the system generating fMRI signals, $u_{i}(t)$ represents the input stimulus, $\boldsymbol{w}_{\boldsymbol{v}}^{(i)}$ refers to the stimulus distribution of signal intensity, and $H(u_{i}(t))$ is the neuronal activation pattern. Therefore, theoretically, $\boldsymbol{x}_{\boldsymbol{v}}$ can be decomposed into the weighted superposition of several time series which characterize neuronal activation patterns. If we can model the neuronal activation patterns and their distributions in fMRI which are shown in Fig. 4, the network can eliminate the huge redundancies in fMRI data.
3.4 Modeling of Neuronal Activation Pattern
Inspired by the above prior, as shown in Fig. 2, we have established a learnable matrix $C_{\text{train}}\in\mathbb{R}^{K\times T}$. Each row of this learnable matrix represents a kind of reusable neuronal activation pattern, and it is mapped to the feature space through Conv1d blocks. As such, the network is capable of learning $K$ types of reusable patterns to model the neuronal activation correlations between all $\boldsymbol{x}_{\boldsymbol{v}}$.
@@ 3.5 Modeling of Activation Pattern Distribution and Feature Fusion
While modeling several reusable neuronal activation patterns, we use INR groups to model their spatial distributions to further decrease the redundancies among each distribution
$g_{i}(\boldsymbol{v}|\sigma_{i})\rightarrow\boldsymbol{w}_{\boldsymbol{v}}^{(i)},i\in[0,K).$ (8)
Here function $g_{i}$ represents the $i$th INR network with $\sigma_{i}$ being its parameters, which models the distribution of the $i$th neuronal activation pattern. For every spatial coordinate $\boldsymbol{v}$, we concatenate the outputs of these $K$ INR networks into
$\boldsymbol{w}_{\boldsymbol{v}}=[\boldsymbol{w}_{\boldsymbol{v}}^{(0)},\boldsymbol{w}_{\boldsymbol{v}}^{(1)},\cdots,\boldsymbol{w}_{\boldsymbol{v}}^{(K-1)}].$ (9)
The components of $\boldsymbol{w}_{\boldsymbol{v}}\in\mathbb{R}^{K}$ represent the spatial distribution values in coordinate $\boldsymbol{v}$ of each activation pattern. In this way, we can further eliminate the redundancies within the distributions of neuronal activities.
@@ 3.6 Activation Pattern Initialization by ICA
In order to improve the convergence process of the neural network, it is crucial to design a suitable initialization for the modeling of neuronal activation patterns, i.d. matrix $C_{train}$. As described by the principle of localization[32] mentioned earlier, the brain can be divided into regions, where are significant correlations among the neuronal activation patterns.
Taking inspiration from this, we use one kind of neuronal activation pattern to describe the main neuronal activation in each certain region. Specifically, we adopted the ICA algorithm to decompose the original fMRI data into basis time series as the neuronal activation patterns, by which we initialize the learnable matrix $C_{train}$. The matrix form of ICA can be represented as $Y=AS$, where $Y$ denotes the original signal matrix decomposed into a mixing matrix $A$ and a source matrix $S=\{\vec{s}_{1},\vec{s}_{2},\cdots\}$. Correspondingly, applying ICA to fMRI signals gets
$\boldsymbol{x}_{\boldsymbol{v}}=\boldsymbol{s}_{\boldsymbol{v}}\cdot\left[\boldsymbol{a}_{0},\boldsymbol{a}_{1},\cdots,\boldsymbol{a}_{K-1}\right]^{T}.$ (10)
In this equation, $\boldsymbol{x}_{\boldsymbol{v}}$ represents the fMRI time series, $\boldsymbol{a}_{i}\in\mathbb{R}^{T\times 1}$ denotes the activation patterns, with the vector $\boldsymbol{s}_{\boldsymbol{v}}=[\boldsymbol{s}_{\boldsymbol{v}}^{(0)},\boldsymbol{s}_{\boldsymbol{v}}^{(1)},\cdots,\boldsymbol{s}_{\boldsymbol{v}}^{(K-1)}]$ describing their spatial distributions. ICA makes the independence between each distributions maximized, which allows the activation patterns we get are the main pattern in each certain region. This method make sure we can better preserve the dominant information, fulfilling the requirement of fidelity for medical data.
Figure 3: An example of neuronal activation patterns and their distributions.
Figure 3: An example of neuronal activation patterns and their distributions.
@@
3.7 Loss Function
We employ a weighted combination of L2 loss and SSIM loss as the network’s loss function, which is formulated as
$l=\left(1-SSIM\left(\{{\boldsymbol{x}_{\boldsymbol{v}}}\}_{\boldsymbol{v}=\boldsymbol{a}}^{\boldsymbol{v}=\boldsymbol{b}},\{\widetilde{\boldsymbol{x}_{\boldsymbol{v}}}\}_{\boldsymbol{v}=\boldsymbol{a}}^{\boldsymbol{v}=\boldsymbol{b}}\right)\right)*\sigma+L_{2}\left(\{{\boldsymbol{x}_{\boldsymbol{v}}}\}_{\boldsymbol{v}=\boldsymbol{a}}^{\boldsymbol{v}=\boldsymbol{b}},\{\widetilde{\boldsymbol{x}_{\boldsymbol{v}}}\}_{\boldsymbol{v}=\boldsymbol{a}}^{\boldsymbol{v}=\boldsymbol{b}}\right)$ (11)
The $\{{\boldsymbol{x}_{\boldsymbol{v}}}\}_{\boldsymbol{v}=\boldsymbol{a}}^{\boldsymbol{v}=\boldsymbol{b}}$ is a batch of ground truth data, and $\{\widetilde{\boldsymbol{x}_{\boldsymbol{v}}}\}_{\boldsymbol{v}=\boldsymbol{a}}^{\boldsymbol{v}=\boldsymbol{b}}$ is the prediction of this batch of data given by the network. $\sigma$ is a parameter to adjust L2 loss and SSIM loss at the same magnitude.